An organization determines that the probability of unauthorized access to a database that contains personally identifiable information (PII) about its clients and employees is 5% in a year. The total estimate of the loss due to this exposure is estimated to be 5 million dollars. This includes losses resulting from loss of reputation, business operations, fines imposed by FCC, legal fees.

After consulting with a security firm, a product was identified that could implement stronger access control and that could allow security administrator to track such an unauthorized access. The total cost of the product is $500,000, plus $10,000/year for maintenance. The product should work well for next five years. It is estimated that, if implemented, it will protect from 90% of all such attacks.

We want to know whether the organization should purchase this product.



What is the Single Loss Expectancy (SLE)? $ (Please enter digits only)

What is the Annualized Rate of Occurrence (ARO)?   %

What is the calculated Annualized Loss Expectancy (ALE)? $  (Please enter digits only)

What would be the ARO if the countermeasure is implemented?  %

What would be the corresponding ALE? $  (Please enter digits only)

What is the countermeasure cost for a year? $  (Please enter digits only)

Should the organization purchase this product?  (Please enter Yes or No)
Answer #1

Asset Value (AV) (assumed)=510,000 USD

Exposure Factor (EF) = 5%

(1) SLE =AV*EF=(510000*5)/100=25500

Total Loss due to this exposure = 5M USD= 50,00,000

Rate of Occurrence (RO in 5 years)=50,00,000/25500=196.5

(2) Annualized Rate of Occurrence=((197/5)/197)*100=20% 0r 40

(3) Annualized Loss Expectancy (ALE)=SLE*ARO= (25500)=25,000*40=10,00.000

After counter measures 90% will not happen so (197*10)/100=20

(4) ARO after counter measure= 20/5=4 or (4/20)*100= 20%

(5) ALE after counter measures= 25500*4=1,02,000

(6) Cost of counter measure is 5,00,000*(5*10,000)=5,50,000 USD for five years.

(7) Yes, The organization should buy this product as loss is reduced.